a line goes through (5,-7) and has a slope m= -3. write the equation that represents the line.
Answer by Cromlix(4381)

Hi there,
Using line equation:
y - b = m(x - a)
m = -3 and (a,b) = (5, -7)
y - (-7) = -3(x - 5)
y + 7 = -3x + 15
y = -3x + 15 - 7
y = -3x + 8
Hope this helps :-)